OpenStudy (anonymous):

Radio signals travel at a rate 3 × 108 of meters per second. How many seconds would it take for a radio signal to travel from a satellite to the surface of the Earth if the satellite is orbiting at a height of 3.6 × 107 meters

OpenStudy (anonymous):

I assume you mean 3*10^8 and 3.6*10^7

OpenStudy (anonymous):

Rate*time=distance time=distance/rate

OpenStudy (anonymous):

yea. i forgot to put the sign.. im pretty sure the answer is 1.2 * 10^-1 but i need the work.

OpenStudy (anonymous):

what did you have for a final answer before you used calculator?

OpenStudy (anonymous):

i jus remembered it, i dont remember how to do it

OpenStudy (anonymous):

yes, that sound correct.

OpenStudy (anonymous):

can show me how to do the work?

OpenStudy (anonymous):

rate=3*10^8m/s distance=3.6*10^7m time=distance/rate=3.6/30

OpenStudy (anonymous):

thank you nsutton